OpenStudy (anonymous):

The number of gallons, P(t), of a pollutant in a lake changes at the rate P'(t) = 1-3e^(-.2sqrt(x)) gallons per day, where t is measured in days. There are 50 gallons of the pollutant in the lake at time t = 0. The lake is considered to be safe when it contains 40 gallons or less of pollutant. a.) Is the amount of pollutant increasing at time t = 9 ? Why or why not? b.) For what value of t will the number of gallons of pollutant be at its minimum? Justify your answer. c.) Is the lake safe when the number of gallons of pollutant is at its minimum? Justify your answer. d.) An investigator..

OpenStudy (anonymous):

d.) An investigator uses the tangent line approximation to P(t) at t = 0 as a model for the amount of pollutant in the lake. At what time t does this model predict that the lake becomes safe?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

a) solve P'(9) if positive, then it is increasing, and vice versa b) set P'(t) = 0 and solve for t

OpenStudy (anonymous):

@epicfail for b.) how do i solve for t with the equation P'(t) = 1- 3e^(-.2sqrt(t))

OpenStudy (anonymous):

well P'(t) = 0 =1-3e^(-2sqrtt)

OpenStudy (anonymous):

^-.2

OpenStudy (anonymous):

so 1=3e^(-2sqrtt)

OpenStudy (anonymous):

you forgot the decimal xD

OpenStudy (anonymous):

1/3=e^-2sqrtt

OpenStudy (anonymous):

ln 1/3= -.2sqrtt

OpenStudy (anonymous):

5ln1/3=-sqrtt

OpenStudy (anonymous):

ln(1/3^5) squared = t

OpenStudy (anonymous):

divide through by 0.2

OpenStudy (anonymous):

nvm :)

OpenStudy (anonymous):

when you take 0.2 to the other side, 0.2=1/5, so effectively you multiply through by 5

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

why did you put down 3^5? @epicfail

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@shamim

OpenStudy (anonymous):

5ln(1/3) = ln((1/3)^5) =ln((1^5)/(3^5)) = ln(1/3^5)

OpenStudy (anonymous):

ok ty :)

OpenStudy (anonymous):

btw if we weren't going to simply .2 to 1/5 can u show me how we do the problem by just using .2 instead of 1/5?

OpenStudy (anonymous):

@epicfail

OpenStudy (anonymous):

exactly the same, ln(...)/0.2=5ln(...)

OpenStudy (anonymous):

okay thank you lol